OpenStudy (anonymous):

what is the simplified form of 18q^4r/6q3r5

OpenStudy (lgbasallote):

\[\huge \implies \frac{18}{6} \times \frac{q^4}{q^3} \times \frac{r}{r^5}\] does that help?

OpenStudy (anonymous):

a.) 3r^4/q B.)12 r^4/q c.)3q/r4 d.)12q/r4

OpenStudy (precal):

 use this law of exponent after lgbasallote's setup